OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

HanAkoSolo (jamierox4ev3r):

To solve questions like this, just make sure you use this formula: \(\Huge\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

y)y-1) x)x-1)

HanAkoSolo (jamierox4ev3r):

well to plug in the points, you would use the y and x values, like so: \(\Huge\frac{-3-3}{-2-1}\)

HanAkoSolo (jamierox4ev3r):

then if you simplify this, you get a slope. What is the slope @wildfire95 ?

HanAkoSolo (jamierox4ev3r):

remember, this formula is only for calculating slope. -3-3=-6, and -2-1=-3 basically you have \(\Huge\frac{-6}{-3}\)

HanAkoSolo (jamierox4ev3r):

simplify that down further and you get a slope of 2x. Now using a point, we will find the y-intercept

OpenStudy (anonymous):

hope this is right y=-2x=1